# A new framework for nonlinear Kalman filters

### v10

@@ 2.2 The conventional framework and its problem
On the other hand, the actual value of the updated state covariance matrix ${\color[rgb]{1,0,0}P_{k|k,ac}}$ given the selected Kalman gain in (9) can be calculated by substituting (9) into (6). Namely:
${\color[rgb]{1,0,0}P_{k|k,ac}}=P_{k|k-1}+{\color[rgb]{1,0,0}P_{xy,k|k-1}}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}S_{k}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}P_{xy,k|k-1}^{T}}$ (11)
$\quad-P_{xy,k}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}P_{xy,k|k-1}^{T}}-{\color[rgb]{1,0,0}P_{xy,k|k-1}}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}P_{xy,k}^{T}$
In other words, ${\color[rgb]{1,0,0}P_{k|k,ac}}$ plays the role of $P_{k|k}$ in (6), and ${\color[rgb]{1,0,0}P_{k|k,est}}$ plays the role of $P_{k|k,op}$ in (8). Intuitively, tr$({\color[rgb]{1,0,0}P_{k|k,ac}})$ tends to be larger than tr$({\color[rgb]{1,0,0}P_{k|k,est}})$ since tr$(P_{k|k,op})=\min$tr$(P_{k|k})\leq$ tr$(P_{k|k})$. However, this relationship of size is not always preserved after the approximations. For example, according to (10) and (11), when $P_{xy,k}=S_{k}=1$ and ${\color[rgb]{1,0,0}P_{xy,k|k-1}}={\color[rgb]{1,0,0}S_{k|k-1}}=0.9$, we have ${\color[rgb]{1,0,0}P_{k|k,est}}-{\color[rgb]{1,0,0}P_{k|k,ac}}=0.1>0$. Nevertheless, in most cases in practice [17, 18], we do find that tr$({\color[rgb]{1,0,0}P_{k|k,ac}})>$tr$({\color[rgb]{1,0,0}P_{k|k,est}})$, meaning that using (10) will often (although not always) underestimate the actual covariance.
In order to make this statement more rigorous in math, we consider the approximated covariance matrices ${\color[rgb]{1,0,0}P_{y,k|k-1}},{\color[rgb]{1,0,0}S_{k|k-1}}$, and ${\color[rgb]{1,0,0}P_{xy,k|k-1}}$ as noisy versions of $P_{y,k},S_{k}$, and $P_{xy,k}$. Specifically, we assume that ${\color[rgb]{1,0,0}P_{xy,k|k-1}}$ is a random matrix, ${\color[rgb]{1,0,0}P_{y,k|k-1}}$ is a random positive semidefinite matrix, and ${\color[rgb]{1,0,0}S_{k|k-1}}$ is a random positive definite matrix. Note that we mark all the random covariance matrices in red in this section for clarity. Additionally, we assume that the “noises” that arise in the approximations are unbiased, so these random matrices should satisfy:
@@
$\begin{cases}\operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}P_{y,k|k-1}}]=P_{y,k}\\ \operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}S_{k|k-1}}]=S_{k}\\ \operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}P_{xy,k|k-1}}]=P_{xy,k}\end{cases}$ (12)
These additional definitions and assumptions are made so that we can consider the “average” case and give a more mathematically rigorous definition of the overconfidence caused by (10). Namely, we have the following theorem.
@@
Suppose that ${\color[rgb]{1,0,0}P_{xy,k|k-1}}$ and ${\color[rgb]{1,0,0}S_{k|k-1}}$ are random matrices that satisfy (12). ${\color[rgb]{1,0,0}S_{k|k-1}}$ is symmetric and positive definite. Then, $P_{k|k,est}$ calculated from (10) satisfies:
$\operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}P_{k|k,ac}}-{\color[rgb]{1,0,0}P_{k|k,est}}]\succeq 0$ (13)
where $\succeq$” represents positive semidefiniteness, and the equality only holds when ${\color[rgb]{1,0,0}P_{xy,k|k-1}}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}$ is constant.
@@
Let ${\color[rgb]{1,0,0}P}={\color[rgb]{1,0,0}P_{xy,k|k-1}}S_{k}^{-\frac{1}{2}}$, ${\color[rgb]{1,0,0}S}=S_{k}^{-\frac{1}{2}}{\color[rgb]{1,0,0}S_{k|k-1}}S_{k}^{-\frac{1}{2}}$. According to (12),
@@
$\begin{cases}\operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}P}]=\bar{P}=P_{xy,k}S_{k}^{-\frac{1}{2}}\\ \operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}S}]=I\end{cases}$ (14)
According to (10, 11),
${\color[rgb]{1,0,0}P_{k|k,ac}}-{\color[rgb]{1,0,0}P_{k|k,est}}$ (15)
$=$ ${\color[rgb]{1,0,0}P}{\color[rgb]{1,0,0}S^{-2}P^{T}}-\bar{P}{\color[rgb]{1,0,0}S^{-1}P^{T}}-{\color[rgb]{1,0,0}PS^{-1}}\bar{P}^{T}+{\color[rgb]{1,0,0}PS^{-1}P^{T}}$
$=$ $({\color[rgb]{1,0,0}PS^{-1}}-\bar{P})({\color[rgb]{1,0,0}PS^{-1}}-\bar{P})^{T}+{\color[rgb]{1,0,0}PS^{-1}P^{T}}-\bar{P}\bar{P}^{T}$
@@
For any fixed unit vector $\bm{v}$, ${\color[rgb]{1,0,0}P^{T}}\bm{v}$ is a random vector with a mean of $\bar{P}^{T}\bm{v}$. According to Lemma 4 in Appendix A,
$\mathbb{E}[\bm{v}^{T}{\color[rgb]{1,0,0}P}{\color[rgb]{1,0,0}S^{-1}}{\color[rgb]{1,0,0}P^{T}}\bm{v}]\geq\bm{v}^{T}\bar{P}\bar{P}^{T}\bm{v}.$ (16)
Moving the right hand side to the left, the following must hold true:
$\mathbb{E}[{\color[rgb]{1,0,0}PS^{-1}P^{T}}-\bar{P}\bar{P}^{T}]\succeq 0$ (17)
@@
Therefore, the expectation of (15) is always p.s.d. due to (17) and the fact that the first term in (15) is p.s.d. The equality only holds when ${\color[rgb]{1,0,0}PS^{-1}}$ is constant, which is equivalent to ${\color[rgb]{1,0,0}K_{est}}={\color[rgb]{1,0,0}P_{xy,k|k-1}}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}$ being constant (i.e., free from approximation errors).
@@
Theorem 1 suggests that using (10) to estimate the state covariance matrix tends to lead to overconfident covariance estimation when the estimated $S_{k}$ and $P_{xy,k}$ are not the same as their actual values. Admittedly, in practice, the errors of the approximated covariance matrices ${\color[rgb]{1,0,0}P_{y,k|k-1}},{\color[rgb]{1,0,0}S_{k|k-1}}$, and ${\color[rgb]{1,0,0}P_{xy,k|k-1}}$ are not as simple as zero-mean random matrices. However, Theorem 1 does give us a structured explanation as to why nonlinear KFs often underestimate the state covariance matrix. Note that an overconfident state covariance estimation can be very detrimental since it prevents the filter from correcting itself in subsequent iterations, thereby affecting the state estimation accuracy. Again, this problem is unavoidable because the actual values of $S_{k}$ and $P_{xy,k}$ are related to the actual state, which is unknown. Therefore, this problem exists for all types of nonlinear KFs as long as (10) is used to update the state covariance estimation.
@@
In plain language, the problem of the conventional framework originates from the assumption that the selected Kalman gain is optimal. However, the Kalman gain, which was optimized based on the approximated measurement function near the predicted states, generally will not minimize the trace of the covariance matrix after the update. The reason is that the states will move to a new location after the update, and the previous approximations may be inaccurate. An example of such a problem is shown in Fig. 1, where we used an EKF to estimate a plane’s position. In the example, the state is the plane’s position along the x-axis, and the measurement function is the altitude of the ground beneath the plane, which is a nonlinear function of the state. Due to the inaccurate approximation of the measurement function around the predicted state (the dashed purple line, “Approx. 1”, in Fig. 1), the EKF overestimates the position of the plane, and the absolute value of the state estimation error $|\hat{x}_{k|k}-x_{k}|$ becomes higher after the “update” step. Meanwhile, the estimated standard deviation of the state, $\sigma_{k|k}$, becomes smaller after the “update” step. Such a contradiction suggests that the effectiveness of the Kalman gain is overestimated. This is visualized by the narrow yellow distribution in the “update” step in Fig. 1, which does not cover the true state.
@@ 2.3 The proposed new framework for nonlinear Kalman filters
Once we fundamentally understand the problem of the conventional framework, the solution is simple. We can redo the approximation after the update and use the newly approximated measurement function to update the covariance estimation. In our new framework, this new step is called “recalibrate”. As shown in Fig. 1, with this additional step, the covariance matrix may not have the smallest possible trace after the update; instead, it will honestly reflect the effect of the Kalman gain on the state covariance matrix. This is visualized by the wider yellow distribution in the “recalibrate” step, which now covers the true state.
@@
Mathematically, “recalibrate” means to re-approximate the measurement functions around the updated states $\hat{x}_{k|k}$ and re-estimate the matrices $P_{y,k},S_{k}$ and $P_{xy}$. We denote these re-approximated covariance matrices as ${\color[rgb]{1,0,0}P_{y,k|k}},{\color[rgb]{1,0,0}S_{k|k}}$, and ${\color[rgb]{1,0,0}P_{xy,k|k}}$ since they are all estimated based on the updated states. In the new framework, the state covariance matrix is updated by replacing $S_{k}$ with ${\color[rgb]{1,0,0}S_{k|k}}$ and $P_{xy,k}$ with ${\color[rgb]{1,0,0}P_{xy,k|k}}$ in (11):
${\color[rgb]{1,0,0}P_{k|k,new}}=P_{k|k-1}+{\color[rgb]{1,0,0}K_{est}}{\color[rgb]{1,0,0}S_{k|k}}{\color[rgb]{1,0,0}K_{est}^{T}}$ (18)
$\quad-{\color[rgb]{1,0,0}P_{xy,k|k}}{\color[rgb]{1,0,0}K_{est}^{T}}-{\color[rgb]{1,0,0}K_{est}}{\color[rgb]{1,0,0}P_{xy,k|k}^{T}}$
$=P_{k|k-1}+{\color[rgb]{1,0,0}P_{xy,k|k-1}}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}S_{k|k}}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}P_{xy,k|k-1}^{T}}$
$\quad-{\color[rgb]{1,0,0}P_{xy,k|k}}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}P_{xy,k|k-1}^{T}}-{\color[rgb]{1,0,0}P_{xy,k|k-1}}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}P_{xy,k|k}^{T}}$

### v11

@@ 2.2 The conventional framework and its problem
On the other hand, the actual value of the updated state covariance matrix $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k,\text{ac}}$} given the selected Kalman gain in (12) can be calculated by substituting (12) into (7). Namely,
${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k,\text{ac}}}=P_{k|k-1}+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}^{-1}}S_{k}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}^{T}}$ (14)
$\quad-P_{xz,k}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}^{T}}-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}^{-1}}P_{xz,k}^{T}.$
The covariance estimation error after the state update can therefore be calculated as
${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k,\mathrm{err}}}:={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k,\text{ac}}}-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k,\text{est}}}.$ (15)
Noticing that the covariance estimation error before the state update is
${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k-1,\mathrm{err}}}:=P_{k|k-1}-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k-1,\text{est}}},$ (16)
the change in the covariance estimation error is
${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\Delta P_{k|k,\mathrm{err}}}={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k,\mathrm{err}}}-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k-1,\mathrm{err}}}$ (17)
$\quad={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k,\mathrm{ac}}}-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k,\mathrm{est}}}|_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k-1,\text{est}}}=P_{k|k-1}}.$
[…]
To formalize this idea, we interpret the approximated covariance matrices $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{z,k|k-1}$}, $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}$}, and $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}$} produced by a nonlinear KF as (random) estimators of the actual quantities $P_{z,k}$, $S_{k}$, and $P_{xz,k}$. More precisely, we regard $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}$} as a random matrix, $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{z,k|k-1}$} as a random positive semidefinite matrix, and $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}$} as a random positive definite matrix. For clarity, all such random covariance matrices are marked in red in this section. Additionally, we assume that these approximated covariance matrices are unbiased estimators and satisfy
@@
$\begin{cases}\operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{z,k|k-1}}]=P_{z,k},\\ \operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}}]=S_{k},\\ \operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}}]=P_{xz,k}.\end{cases}$ (18)
These additional definitions and assumptions are introduced to facilitate consideration of the “average” case and provide a more mathematically rigorous definition of the overconfidence caused by (13). Specifically, (18) should be understood as follows. If a systematic, nonzero-mean bias in these approximations were known a priori, it would be natural to modify the filter so as to subtract this bias, and to analyze only the residual approximation error. Thus, it is reasonable to model the remaining covariance errors as zero-mean random matrices. Note that this viewpoint is conceptually analogous to the standard treatment of deterministic biases in augmented-state or separate-bias Kalman filters, where unknown biases are initialized as zero-mean random variables. Namely, any known mean component is absorbed into the system model, and only the zero-mean uncertainty is left for the filter to estimate.
With (18), we have the following theorem about the overconfidence of nonlinear Kalman filters.
@@
Suppose that $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}$} and $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}$} are random matrices that satisfy (18). $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}$} is symmetric and positive definite. Then, $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\Delta P_{k|k,\mathrm{err}}$} calculated from (17) satisfies:
$\operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\Delta P_{k|k,\mathrm{err}}}]\succeq 0,$ (19)
where $\succeq$” represents positive semidefiniteness, and the equality only holds when ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}^{-1}}$ is constant almost surely. Note that this theorem does not require $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}$} and $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}$} to be independent.
@@
Let ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P}={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}}S_{k}^{-\frac{1}{2}}$, ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S}=S_{k}^{-\frac{1}{2}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}}S_{k}^{-\frac{1}{2}}$. According to (18),
@@
$\begin{cases}\operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P}]=\bar{P}=P_{xz,k}S_{k}^{-\frac{1}{2}},\\ \operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S}]=I.\end{cases}$ (20)
According to (13, 14, 18),
${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\Delta P_{k|k,\mathrm{err}}}$ (21)
$=$ ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S^{-2}P^{T}}-\bar{P}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S^{-1}P^{T}}-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}PS^{-1}}\bar{P}^{T}+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}PS^{-1}P^{T}}$
$=$ $({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}PS^{-1}}-\bar{P})({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}PS^{-1}}-\bar{P})^{T}+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}PS^{-1}P^{T}}-\bar{P}\bar{P}^{T}.$
@@
For any fixed unit vector $\bm{v}$, ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P^{T}}\bm{v}$ is a random vector with a mean of $\bar{P}^{T}\bm{v}$. According to Lemma 4 in Appendix A,
$\mathbb{E}[\bm{v}^{T}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S^{-1}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P^{T}}\bm{v}]\geq\bm{v}^{T}\bar{P}\bar{P}^{T}\bm{v}.$ (22)
Moving the right hand side to the left, we get
$\mathbb{E}[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}PS^{-1}P^{T}}-\bar{P}\bar{P}^{T}]\succeq 0.$ (23)
@@
Therefore, the expectation of (21) is always positive semi-definite. due to (23) and the fact that the first term in (21) is positive semi-definite. The equality only holds when $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}PS^{-1}$} is constant almost surely, which is equivalent to ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}K_{k,\text{est}}}={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}^{-1}}$ being constant (i.e., free from approximation errors) almost surely.
@@
Theorem 1 suggests that using (13) to estimate the state covariance matrix tends to lead to overconfident covariance estimation when the estimated $S_{k}$ and $P_{xz,k}$ are not the same as their actual values. Admittedly, in practice, the errors of the approximated covariance matrices ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{z,k|k-1}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}}$, and $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}$} may not be as simple as zero-mean random matrices. However, Theorem 1 does give us a structured explanation as to why nonlinear KFs often underestimate the state covariance matrix. Note that an overconfident state covariance estimation can be very detrimental since it prevents the filter from correcting itself in subsequent iterations, thereby affecting the state estimation accuracy. Again, this problem is unavoidable because the actual values of $S_{k}$ and $P_{xz,k}$ are related to the actual distribution of the states, which is unknown. Therefore, this problem exists for all types of nonlinear KFs as long as (13) is used to update the state covariance estimation.
@@
In plain language, the problem of the conventional framework originates from the assumption that the selected Kalman gain is optimal. However, the Kalman gain, which was optimized based on the approximated measurement function near the predicted states, generally will not minimize the trace of the actual covariance. The reason is that the actual states can be far from the predicted states and the approximations made in the “predict” step may be inaccurate. An example of such a problem is shown in Fig. 1, where we use an EKF to estimate a plane’s position. In the example, the state is the plane’s position along the x-axis, and the measurement function is the altitude of the ground beneath the plane, which is a nonlinear function of the state. Due to the inaccurate approximation of the measurement function around the predicted state (the dashed purple line, “Approx. 1”, in Fig. 1), the EKF overestimates the position of the plane, and the absolute value of the state estimation error $|x_{k}-\hat{x}_{k|k}|$ becomes higher after the “update” step. Meanwhile, the estimated standard deviation of the state, $\sigma_{k|k}$, becomes smaller after the “update” step. Such a contradiction suggests that the effectiveness of the Kalman gain is overestimated. This is visualized by the narrow yellow distribution in the “update” step in Fig. 1, which does not cover the true state.
@@ 2.3 The proposed new framework for nonlinear Kalman filters
Once we fundamentally understand the problem of the conventional framework, the solution is simple. We can redo the approximation after the update and use the newly approximated measurement function to update the covariance estimation. In our new framework, this new step is called “recalibrate”. As shown in Fig. 1, with this additional step, the estimated covariance matrix may not have the smallest possible trace after the update; instead, it will honestly reflect the effect of the Kalman gain on the actual state covariance matrix. This is visualized by the wider yellow distribution in the “recalibrate” step, which now covers the true state.
@@
Mathematically, “recalibrate” means to re-approximate the measurement functions around the updated states $\hat{x}_{k|k}$ and re-estimate the matrices $P_{z,k},S_{k}$ and $P_{xz}$. We denote these re-approximated covariance matrices as ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{z,k|k}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k}}$, and $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k}$} since they are all estimated based on the updated states. Similar to (11), these covariance matrices are calculated by
$\begin{cases}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{z,k|k}}:=\Phi_{y}(h;\hat{x}_{k|k},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k-1,\text{est}}}),\\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k}}:={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{z,k|k}}+R_{k},\\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k}}:=\Phi_{xy}(h;\hat{x}_{k|k},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k-1,\text{est}}}).\end{cases}$ (24)
In the new framework, the state covariance matrix is updated by replacing $S_{k}$ with $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k}$} and $P_{xz,k}$ with $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k}$} in (14):
${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k,\mathrm{est,new}}}={{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k-1,\text{est}}}}+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}K_{k,\text{est}}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}K_{\text{est}}^{T}}$ (25)
$\quad-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}K_{\text{est}}^{T}}-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}K_{k,\text{est}}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k}^{T}}$
$={{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k-1,\text{est}}}}+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}^{T}}$
$\quad-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}^{T}}-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k}^{T}}.$
In the new framework, the covariance estimation error after the state update can then be calculated as
${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k,\mathrm{err,new}}}:={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k,\text{ac}}}-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k,\mathrm{est,new}}}.$ (26)
The change in the covariance estimation error after the update is
${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\Delta P_{k|k,\mathrm{err,new}}}={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k,\mathrm{err}}}-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k-1,\mathrm{err,new}}}$ (27)
$\quad={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k,\mathrm{ac}}}-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k,\mathrm{est,new}}}|_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k-1,\text{est}}}=P_{k|k-1}}.$